x=6-4y
2x-3y=1
i dont know how to do these and i have 20 missing assignments please help!

Answer by Maths68(1474)

x=6-4y........(1)
2x-3y=1........(2)
Put the value of x in (2)
2(6-4y)-3y=1
12-8y-3y=1
-11y=1-12
-11y=-11
Divide by -11
-11y/-11=-11/-11
y=1
Put the value of y in (1)
x=6-4y
x=6-4(1)
x=6-4
x=2
Solution Set
x,y=2,1

Check
=====
Put the values of x and y in (1)
x=6-4y........(1)
2=6-4(1)
2=6-4
2=2
Put the values of x and y in (2)
2x-3y=1........(2)
2(2)-3(1)=1
4-3=1
1=1